The sum of Richards age and his older brothers age is 63. The difference between their age is 13. How old are Richard and his brother? Explain how you found your answer.
Answer by ankor@dixie-net.com(22740)

let r = Richard's present age
let b = his brothers
:
Write an equation for each statement
"The sum of Richards age and his older brothers age is 63."
r + b = 63
" The difference between their age is 13."
r - b = 13
How old are Richard and his brother?
We can use elimination with these two equations
r + b = 63
r - b = 13
------------Adding these two equations eliminate b, find r
2r + 0 = 76
r = 76/2
r = 38 yrs is Richard's age
:
Find the b using the 1st equation, replace r with 38
38 + b = 63
b = 63 - 38
b = 25 yrs is brother's age
:
:
:
Check solutions in the 2nd equation
38 - 25 = 13